what is the y intercept from the following example. triangle RST has a scale factor of 1/2 and coordinates of the following r( 0, 2) s(6, -4) T( -4,-12) i keep getting -22/5 and the real answer is -17/5 with a slope of 4/5. the slope is no problem but how are they getting -17/5 ????
Found 2 solutions by drk, MRperkins:
Answer by drk(1908)

We know that
R(0,2), S(6,-4), and T (-4,-12).
Applying a scale factor of 1/2, we get
R'(0,1), S'(3,-2), and T'(-2,-6).
The slope of ST = slope of S'T' = 4/5.
The Y-intercept of S'T' = -22/5.
The Y-intercept of ST = -44/5.
I don't see anything wrong with your answer. Is it possible that the "real" answer or book answer is incorrect?

Answer by MRperkins(300)

I have worked this problem as well and I, too, get the y-intercept of -22/5.
I just wanted to reinforce this.